### v2

@@ Abstract
Large vision-language models (LVLMs) excel at multimodal understanding but suffer from high computational costs due to redundant vision tokens. Existing pruning methods typically rely on single-layer attention scores to rank and prune redundant visual tokens to solve this inefficiency. However, as the interaction between tokens and layers is complicated, this raises a basic question: Is such a simple single-layer criterion sufficient to identify redundancy? To answer this question, we rethink the emergence of redundant visual tokens from a fundamental perspective: information flow, which models the interaction between tokens and layers by capturing how information moves between tokens across layers. We find (1) the CLS token acts as an information relay, which can simplify the complicated flow analysis; (2) the redundancy emerges progressively and dynamically via layer-wise attention concentration; and (3) relying solely on attention scores from single layers can lead to contradictory redundancy identification. Based on this, we propose FlowCut, an information-flow-aware pruning framework, mitigating the insufficiency of the current criterion for identifying redundant tokens and better aligning with the model’s inherent behaviors. Extensive experiments show that FlowCut achieves superior results, outperforming SoTA by 1.6% on LLaVA-1.5-7B with 88.9% token reduction, and by 4.3% on LLaVA-NeXT-7B with 94.4% reduction, delivering 3.2$\times$ speed-up in the prefilling stage. Our code is available at https://github.com/TungChintao/FlowCut.
@@ 1 Introduction
    We provide a novel perspective (information flow) for understanding the emergence of redundant tokens, revealing redundancy as a natural consequence of asymmetric attention.
@@ 3.2 Multi-Criteria Evaluator
$\mathbf{I}^{\textrm{a}}=\mathbf{A}^{\textrm{g}},\ \ \mathbf{I}^{\textrm{s}}=\mathrm{Softmax}\left(\frac{\mathbf{V_{g}}\mathbf{V}^{\top}}{\sqrt{D_{v}}}\right),\ \ \mathbf{I}^{\textrm{r}}=\left\lVert\mathbf{V}\right\rVert_{1}$ (4)
@@ 3.3 Cumulative Flow Importance Tracking
Table 1: Performance of FlowCut on LLaVA-1.5 under different visual token settings. The vanilla number of visual tokens is 576. The last column is the average accuracy proportion relative to the upper limit. FlowCut achieves optimal performance on various image understanding benchmarks.
@@ 4.1 Effectiveness of FlowCut in Inference
Our proposed FlowCut substantially improves inference efficiency in LVLMs while maintaining comparable performance. As shown in Table 5, we compare total inference time, prefilling time, and FLOPs on both LLaVA-1.5-7B and LLaVA-NeXT-7B against several existing methods. FlowCut consistently achieves the best efficiency gains. Notably, it delivers a 3.2$\times$ speedup in prefilling and a 3.0$\times$ speedup in the entire inference on LLaVA-NeXT.
@@
LLaVA-1.5-7B 576 17:05 1.0$\times$ 102ms 1.0$\times$ 8.82
+ FastV 64 15:32 1.1$\times$ 87.3ms 1.2$\times$ 2.26
+ SparseVLM 64 15:57 1.1$\times$ 90.1ms 1.1$\times$ 2.31
+ PDrop 64 12:56 1.3$\times$ 72.5ms 1.4$\times$ 2.16
+ VisionZip 64 12:10 1.4$\times$ 69.2ms 1.5$\times$ 2.03
+ FlowCut 64 11:17 1.5$\times$ 62.6ms 1.6$\times$ 1.95
@@
LLaVA-NeXT-7B 2880 35:16 1.0$\times$ 206ms 1.0$\times$ 31.03
+ FastV 160 28:27  1.3$\times$ 117ms 1.8$\times$ 7.35
+ SparseVLM 160 30:26  1.2$\times$ 136ms 1.5$\times$ 7.62
+ PDrop 160 13:45  2.6$\times$ 79.5ms 2.6$\times$ 6.78
+ VisionZip 160 12:32  2.8$\times$ 69.9ms 2.9$\times$ 4.72
+ FlowCut 160 11:40 3.0$\times$ 63.8ms 3.2$\times$ 4.29
@@ 4.2 Effectiveness of FlowCut in Training
Our method can also be seamlessly integrated into models during training to reduce the number of visual tokens, thereby saving memory and shortening training time. As shown in Table 6, we evaluate FlowCut with different token configurations across 10 benchmarks on LLaVA-1.5-7B. The results demonstrate that with 192 tokens, FlowCut achieves a 1.5$\times$ training speedup while actually improving performance to 101.1% of the original model. Even with further reduction to 128 tokens, training speedup reaches 1.6$\times$ while still retaining 99.8% of the original performance.
@@ A.1 Dataset
GQA [19] The GQA benchmark is structured around three parts: scene graphs, questions, and images. It enriches visual content with comprehensive spatial information and object-level attributes. Questions are specifically designed to evaluate models’ ability to comprehend visual scenes and reason about diverse aspects of images.
@@
MMBench [35] The MMBench Benchmark evaluates models through three hierarchical levels of abilities. The first level (L-1) focuses on fundamental perception and reasoning capabilities. The second level (L-2) expands into six distinct sub-abilities, while the third level (L-3) further refines these into 20 specific dimensions. This structure enables a granular and comprehensive assessment of a model’s various capabilities. Additionally, MMB-CN is the Chinese version of the benchmark.
@@
MME [12] MME comprehensively evaluates models’ perceptual and cognitive abilities across 14 subtasks. By employing manually constructed instruction-answer pairs and concise instructions, it effectively minimizes data leakage and ensures fair assessment of model performance.
@@
POPE [29] The POPE benchmark systematically evaluates object hallucination in models through a series of binary questions about object presence in images. Using accuracy, recall, precision, and F1 score as metrics, it precisely measures hallucination levels across different sampling strategies.
@@
ScienceQA [36] ScienceQA encompasses a wide array of domains, including natural, language, and social sciences, with questions hierarchically organized into 26 topics, 127 categories, and 379 skills. Through this comprehensive structure, it offers diverse scientific questions that effectively evaluate multimodal understanding, multi-step reasoning capabilities, and interpretability.
@@
VQA-V2 [14] VQA-V2 evaluates models’ visual perception capabilities through open-ended questions about 265,016 real-world scence images. Each question contains 10 human-annotated ground truth answers, enabling thorough assessment of a model’s ability to interpret and respond to visual queries.
@@
TextVQA [39] TextVQA focuses on the integration of textual information within images. It evaluates a model’s ability to interpret visual elements and embedded text in images through tasks, requiring both visual and textual comprehension to answer questions accurately.
@@
VizWiz [17] The VizWiz benchmark contains over 31,000 visual questions created by blind individuals who photographed scenes with mobile phones while recording spoken questions. Each question includes 10 crowdsourced answers. The dataset presents distinct challenges: lower-quality images from visually impaired photographers, conversational spoken questions, and some questions that remain unanswerable due to content limitations.
@@
SEEDBench [24] SEEDBench contains 19,000 human-annotated multiple-choice questions evaluating models across 12 distinct aspects. It comprehensively assesses capabilities in recognizing spatial and temporal patterns within both images and videos.
@@
MMVet [49] MMVet defines six core abilities (recognition, OCR, knowledge, language generation, spatial awareness, and math) and evaluates models through 16 specific capability combinations. These integrations enable comprehensive assessment of models’ performance across complex scenarios.
@@
LLaVA-Bench [32] LLaVA-Bench categorizes questions into three types: conversational (simple QA), detailed description, and complex reasoning tasks, featuring 24 diverse images with 60 questions spanning indoor and outdoor scenes, memes, paintings, sketches, and more. Each image contains a manually curated detailed description and carefully selected questions, designed to assess model robustness across various prompts.
@@
TGIF-QA [20] TGIF-QA extends the image question answering task to videos with 165,000 question-answer pairs based on GIFs. It introduces four task types: three requiring spatio-temporal reasoning (repetition count, repeating action, state transition) and frame QA answerable from single frames.
@@
MSVD-QA [47] The MSVD-QA benchmark, based on the MSVD dataset, features 1,970 video clips with approximately 50.5K question-answer pairs. It presents open-ended questions across five categories (what, who, how, when, where) covering diverse video content aspects, serving both video question answering and captioning tasks.
@@
MSRVTT-QA [47] MSRVTT-QA comprises 10K video clips with 243K question-answer pairs, challenging models to integrate visual and temporal elements of video content. Similar to MSVD-QA, it includes five question types, evaluating models’ ability to comprehend complex video content.
@@ A.2 Related Work
Recently, building on the success of large language models (LLMs) [2, 3, 8, 43], LVLMs [4, 7, 32, 34, 40, 42] have demonstrated impressive performance in multimodal reasoning by bridging visual and linguistic modalities. However, a critical challenge persists in processing images: the quadratic growth of visual tokens as image resolution increases, which leads to prohibitive computational costs during training and inference. High-resolution image models like LLaVA-NeXT [33] and mini-Gemini-HD [27] process thousands of tokens per image, while video models such as VideoLLaVA [30] and VideoPoet [22] demand even more tokens for multiple frames. This highlights the need for more efficient methods to extract information from visual tokens, instead of merely extending their length.
@@
Efficient inference [10, 9, 23, 31] in LLMs is challenged by their autoregressive nature, where each token prediction depends on the full preceding context. Recently, a variety of token reduction strategies have been proposed to accelerate inference and compress the key-value (KV) cache [18]. For instance, StreamingLLM [45] retains only attention sinks and most recent tokens to shrink the KV cache, while FastGen [13] adaptively manages KV cache based on the behavior of attention head. Heavy-Hitter Oracle (H2O) [51] employs accumulated attention-based scoring to prune key-value pairs during the generation stage. These methods primarily target textual redundancy to improve inference efficiency.
A.3 Broader Impact

### v3

@@ Abstract
[…]
@@ 1 Introduction
    We provide a unified, bottom-up perspective (information flow) by modeling token-to-token interaction to explore the propagation pattern of visual information and understand the emergence of redundant, revealing redundancy as a natural consequence of asymmetric attention.
@@ 3.2 Multi-Criteria Evaluator
$\mathbf{I}^{\textrm{a}}=\mathbf{A}^{\textrm{g}},\ \ \mathbf{I}^{\textrm{s}}=\mathrm{Softmax}\left(\frac{\mathbf{V_{g}}\mathbf{V}^{\top}}{\sqrt{D_{v}}}\right),\ \ \mathbf{I}^{\textrm{d}}=\left\lVert\mathbf{V}\right\rVert_{1}$ (4)
@@ 3.3 Cumulative Flow Importance Tracking
Table 1: Performance of FlowCut on LLaVA-1.5 under different visual token settings. The vanilla number of visual tokens is 576. The last column is the average accuracy proportion relative to the upper limit. FlowCut achieves optimal performance on various image understanding benchmarks. The comparison with additional methods and setting is provided in Appendix A.3.
@@ 4.1 Effectiveness of FlowCut in Inference
[…]
@@
[…]
@@
[…]
@@ 4.2 Effectiveness of FlowCut in Training
[…]
@@ 6 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China under grants 62206102; the National Key Research and Development Program of China under grant 2024YFC3307900; the National Natural Science Foundation of China under grants 62436003, 62376103 and 62302184; Major Science and Technology Project of Hubei Province under grant 2025BAB011 and 2024BAA008; Hubei Science and Technology Talent Service Project under grant 2024DJC078; and Ant Group through CCF-Ant Research Fund. The computation is completed in the HPC Platform of Huazhong University of Science and Technology.
@@ A.1 Dataset
GQA hudson2019gqa The GQA benchmark is structured around three parts: scene graphs, questions, and images. It enriches visual content with comprehensive spatial information and object-level attributes. Questions are specifically designed to evaluate models’ ability to comprehend visual scenes and reason about diverse aspects of images.
@@
MMBench liu2024mmbench The MMBench Benchmark evaluates models through three hierarchical levels of abilities. The first level (L-1) focuses on fundamental perception and reasoning capabilities. The second level (L-2) expands into six distinct sub-abilities, while the third level (L-3) further refines these into 20 specific dimensions. This structure enables a granular and comprehensive assessment of a model’s various capabilities. Additionally, MMB-CN is the Chinese version of the benchmark.
@@
MME fu2023mme MME comprehensively evaluates models’ perceptual and cognitive abilities across 14 subtasks. By employing manually constructed instruction-answer pairs and concise instructions, it effectively minimizes data leakage and ensures fair assessment of model performance.
@@
POPE li2023evaluating The POPE benchmark systematically evaluates object hallucination in models through a series of binary questions about object presence in images. Using accuracy, recall, precision, and F1 score as metrics, it precisely measures hallucination levels across different sampling strategies.
@@
ScienceQA lu2022learn ScienceQA encompasses a wide array of domains, including natural, language, and social sciences, with questions hierarchically organized into 26 topics, 127 categories, and 379 skills. Through this comprehensive structure, it offers diverse scientific questions that effectively evaluate multimodal understanding, multi-step reasoning capabilities, and interpretability.
@@
VQA-V2 goyal2017making VQA-V2 evaluates models’ visual perception capabilities through open-ended questions about 265,016 real-world scence images. Each question contains 10 human-annotated ground truth answers, enabling thorough assessment of a model’s ability to interpret and respond to visual queries.
@@
TextVQA singh2019towards TextVQA focuses on the integration of textual information within images. It evaluates a model’s ability to interpret visual elements and embedded text in images through tasks, requiring both visual and textual comprehension to answer questions accurately.
@@
VizWiz gurari2018vizwiz The VizWiz benchmark contains over 31,000 visual questions created by blind individuals who photographed scenes with mobile phones while recording spoken questions. Each question includes 10 crowdsourced answers. The dataset presents distinct challenges: lower-quality images from visually impaired photographers, conversational spoken questions, and some questions that remain unanswerable due to content limitations.
@@
SEEDBench li2024seed SEEDBench contains 19,000 human-annotated multiple-choice questions evaluating models across 12 distinct aspects. It comprehensively assesses capabilities in recognizing spatial and temporal patterns within both images and videos.
@@
MMVet yu2023mm MMVet defines six core abilities (recognition, OCR, knowledge, language generation, spatial awareness, and math) and evaluates models through 16 specific capability combinations. These integrations enable comprehensive assessment of models’ performance across complex scenarios.
@@
LLaVA-Bench liu2024improved LLaVA-Bench categorizes questions into three types: conversational (simple QA), detailed description, and complex reasoning tasks, featuring 24 diverse images with 60 questions spanning indoor and outdoor scenes, memes, paintings, sketches, and more. Each image contains a manually curated detailed description and carefully selected questions, designed to assess model robustness across various prompts.
@@
TGIF-QA jang2017tgif TGIF-QA extends the image question answering task to videos with 165,000 question-answer pairs based on GIFs. It introduces four task types: three requiring spatio-temporal reasoning (repetition count, repeating action, state transition) and frame QA answerable from single frames.
@@
MSVD-QA xu2017video The MSVD-QA benchmark, based on the MSVD dataset, features 1,970 video clips with approximately 50.5K question-answer pairs. It presents open-ended questions across five categories (what, who, how, when, where) covering diverse video content aspects, serving both video question answering and captioning tasks.
@@
MSRVTT-QA xu2017video MSRVTT-QA comprises 10K video clips with 243K question-answer pairs, challenging models to integrate visual and temporal elements of video content. Similar to MSVD-QA, it includes five question types, evaluating models’ ability to comprehend complex video content.
@@ A.2 Related Work
Recently, building on the success of large language models (LLMs) achiam2023gpt; bai2023qwen; chiang2023vicuna; touvron2023llama, LVLMs bai2023qwenvl; chen2024internvl; liu2024improved; liu2023visual; team2023gemini; tong2024cambrian have demonstrated impressive performance in multimodal reasoning by bridging visual and linguistic modalities. However, a critical challenge persists in processing images: the quadratic growth of visual tokens as image resolution increases, which leads to prohibitive computational costs during training and inference. High-resolution image models like LLaVA-NeXT liu2024llavanext and mini-Gemini-HD li2024mini process thousands of tokens per image, while video models such as VideoLLaVA lin2024video and VideoPoet kondratyuk2024videopoet demand even more tokens for multiple frames. This highlights the need for more efficient methods li2025beyond; li2025bora; li2025toplora; dfn to extract information from visual tokens, instead of merely extending their length.
@@
Efficient inference dao2022flashattention; dao2023flashattention; kwon2023efficient; liu2023ring in LLMs is challenged by their autoregressive nature, where each token prediction depends on the full preceding context. Recently, a variety of token reduction strategies have been proposed to accelerate inference and compress the key-value (KV) cache han2023lm. For instance, StreamingLLM xiao2024efficient retains only attention sinks and most recent tokens to shrink the KV cache, while FastGen ge2023model adaptively manages KV cache based on the behavior of attention head. Heavy-Hitter Oracle (H2O) zhang2023h2o employs accumulated attention-based scoring to prune key-value pairs during the generation stage. These methods primarily target textual redundancy to improve inference efficiency li2024fedbat; li2023alpt.
A.3 Evaluation under Additional Settings and Comparison with More Methods
Retain less visual tokens
To further validate our method under highly constrained token budgets, we conduct experiments with only 32 visual tokens. As shown in Table 8, our method outperforms existing methods by a notable margin under aggressive token reduction, validating its effectiveness in preserving informative content despite severe compression.
Table 8: Performance of FlowCut on LLaVA-1.5 under 32 visual token setting.
Method Tokens GQA TextVQA VQA-V2 POPE SQA VizWiz Avg
LLaVA-1.5 7B (upper limit) 576 61.9 58.2 78.5 85.9 69.5 50.0 100%
FastV 32 46.2 51.5 56.0 35.7 68.3 49.1 78.7%
VisionZip 32 51.5 51.7 65.1 68.3 68.2 49.9 88.7%
FlowCut (Ours) 32 52.1 53.0 66.9 69.6 68.7 52.3 90.8%
Compare with more methods
We have already compared with several representative methods in the main text. Here, we further include more prior works for comparison: LLaVA-PruMerge shang2024llava clusters the visual tokens via k-nearest merging algorithm and merges them. ST3 zhuang2025st3 prunes low-importance tokens progressively across layers using token attention maps. VTW lin2025boosting removes unimportant visual tokens based on attention score at specific layers of the LLM. RCom chen2025recoverable merges unimportant patch tokens based on their similarity to the CLS token and the text tokens. FasterVLM zhang2024cls prunes vision tokens based on CLS attention scores at vision encoder’s output layer.
These methods all rely on single-layer, single-metric scores (e.g., attention or similarity) to assess token importance. In contrast, our method employs multiple metrics and aggregates scores across layers using an EMD-based strategy, which helps mitigate single-metric bias and layer bias zou2021revisiting; tong2025self. We also discuss our differences with CrossGET shi2024crossget here. CrossGET injects learnable cross tokens to capture cross-modal information and merges less important tokens based on cosine similarity. In contrast, our method leverages CLS attention, performs multi-layer and multi-metric evaluation, and applies dynamic layer-wise pruning. Moreover, CrossGET requires re-training due to its learnable cross tokens while our approach is entirely training-free.
As shown in Table 9, our method better preserves performance under similar FLOPs, achieving consistently stronger results across a wide range of benchmarks. Note that for ST3 , as the code is not publicly available, we follow their benchmark protocol for fair comparison.
Table 9: Comparison of FlowCut on LLaVA-1.5-7B with more methods.
[…]
LLaVA-1.5 7B 8.82 58.2 78.5 85.9 64.7 69.5
PruMerge shang2024llava 1.95 53.5 65.9 70.7 56.8 68.5
VTW lin2025boosting (K=5) 2.01 8.1 42.7 46.0 21.3 65.3
RCom chen2025recoverable 1.96 55.5 70.4 72.0 57.9 69.0
FasterVLM zhang2024cls 1.97 55.2 71.9 76.1 60.4 68.9
FlowCut (Ours) 1.95 55.6 72.8 80.2 60.8 69.1
Metric ST3 zhuang2025st3 FlowCut (Ours)
AI2D 55.4 55.8
SQA 68.9 69.2
MMMU 35.3 36.2
MMB 63.8 64.3
POPE 85.2 86.1
TFLOPs 4.27 4.25
A.4 More Sensitivity Analyses of Hyperparameters
Cumulative importance mechanism’s weighting coefficients
We default to a setting of 0.5 for historical scores and 0.5 for current scores. As shown in Table 10, we further conduct a sensitivity analysis. When only the current score is used (i.e., without the cumulative strategy), the POPE performance is 83.8. The best performance is achieved at 0.5:0.5 and 0.6:0.4, while all other settings still outperform the non-cumulative baseline. This demonstrates that our method is robust to changes in the weighting coefficients and validates the effectiveness of the cumulative strategy.
Table 10: Sensitivity analysis of cumulative importance mechanism’s weighting coefficients.
history:current 0:10 (w/o cumulative strategy ) 1:9 2:8 3:7 4:6 5:5 6:4 7:3 8:2 9:1
POPE benchmark 83.8 (baseline) 84.6 84.8 84.6 85.0 85.2 85.2 85.1 85.0 84.7
Pruning frequency
For pruning frequency, we default to pruning every two layers with dynamic prune ratios. As shown in Table 11: 1) Pruning every two layers yields the best performance. Performance decreases as the interval increases, and pruning every layer (n=1) performs slightly worse than n=2, which we attribute to the fact that the first pruning step under n=1 cannot leverage historical scores and must rely solely on the current layer. 2) Multi-layer pruning outperforms single-layer pruning (i.e., only pruning at the last layer), further demonstrating both robustness to this hyperparameter and the benefit of pruning redundancies as they emerge. 3) These results confirm that our method is not sensitive to the prune frequency, highlighting its robustness.
Table 11: Sensitivity analysis of pruning frequency.
Pruning each n layer n=1 n=2 n=3 n=4 n=6 n=8 n=12 only prune at last layer
POPE benchmark 84.9 85.2 84.6 84.6 84.3 84.3 84.0 83.9 (baseline)
A.5 Pseudocode
Here, we provide the pseudocode of our method. And note that method is indeed compatible with FlashAttention because it does not require the full $N\times N$ attention map. Instead, it only needs the $1\times N$ attention value of a single CLS token or global token. When using FlashAttention, we can obtain the necessary attention values with a small, separate computation: 1) For models with a CLS token, we separately compute its $1\times N$ attention map with the patch tokens; 2) For models without a CLS token, we derive a global token (by averaging patch tokens) and then compute its $1\times N$ attention map. This targeted computation is highly efficient, with a complexity of just $O(n)$, adding virtually no overhead. This allows our method and FlashAttention to be used together effectively.
Algorithm 1 Pseudocode for FlowCut
# Input: current layer’s hidden_states and QKV
# Output: selected tokens for current layer
# QKV shape: [B, N, D] (batch size, sequence length, embedding dim)
# If the architecture without CLS token, use a globally pooled token as a substitute
if architecture_without_CLS: cls_token = output.hidden_states.mean(dim=1) # [B,D]
else: cls_token = output.hidden_states[:,0] # [B,D]
patch_token = output.hidden_states[:,1:] # [B,N-1,D]
Q_cls = Q[:, 0] # [B, D], first token is CLS token
K_patch = K[:, 1:] # [B, N-1, D], exclude CLS token
# Compute CLS-to-token attention (O(N) complexity, compatible with FlashAttention)
cls_attn = softmax(dot(Q_cls, K_patch.transpose(-1, -2)) / sqrt(D)) # [B, N-1]
# Compute semantic similarity to evaluate token-level informativeness
V_cls = V[:, 0] # [B, D]
V_patch = V[:, 1:] # [B, N-1, D]
semantic_attn = softmax(dot(V_cls, V_patch.transpose(-1, -2)) / sqrt(D)) # [B, N-1]
# Compute value norm as a proxy for token information capacity
v_score = L1_norm(V_patch) # [B, N-1]
# Fuse multiple importance criteria
importance = (normalize(cls_attn) + normalize(semantic_attn)) * v_score # [B, N-1]
# Accumulate importance across layers (EMA-like smoothing)
cumulative_score = 0.5 * cumulative_score + 0.5 * importance
# Estimate pruning ratio based on CLS attention entropy
entropy = compute_entropy(cls_attn) # [B]
entropy_ratio = entropy / log(N)
prune_ratio = (N - target_num) / sqrt(remain_layer) * (1 - entropy_ratio ** 2)
# Select top-K tokens based on cumulative importance scores
K_keep = round((1 - prune_ratio) * N)
keep_idx = topk(cumulative_score, K=K_keep) # [B, K_keep]
# Retain tokens
if not final_step: tokens = concat(cls_token, filter(patch_tokens, keep_idx))
elif final_step: tokens = filter(patch_tokens, keep_idx)
A.6 Broader Impact